# A Local Projection Stabilised HHO Method for the Oseen Problem

### v1

@@ 2 Discrete Oseen Problem with LPS stabilisation
In this section, we introduce the LPS stabilised discrete formulation for the Oseen problem (1.2) on the hybrid space $\underline{\boldsymbol{U}}_{h,0}^{k}\times P_{h}^{k}$. This section is divided into three sub-sections. The first part defines some reconstruction operators which are essential to define the HHO scheme. In the second part, we discuss a generalised local projection stabilisation setup. The discrete LPS-HHO scheme is defined in the third sub-section.
@@ 2.1 Local Reconstructions
We define three reconstruction operators on the local spaces $\underline{\boldsymbol{U}}_{T}^{k}$, see Oseen_HHO . These will be used to define the discrete HHO bilinear form in (2.7).
Local velocity reconstruction: The velocity reconstruction operator $\boldsymbol{r}_{T}^{k+1}:\underline{\boldsymbol{U}}_{T}^{k}\rightarrow[P^{k+1}(T)]^{2}$ is defined as follows: For any $\underline{\boldsymbol{v}}_{T}\in\underline{\boldsymbol{U}}_{T}^{k}$, $\boldsymbol{r}_{T}^{k+1}\underline{\boldsymbol{v}}_{T}$ must satisfy
$(\nabla(\boldsymbol{r}_{T}^{k+1}\underline{\boldsymbol{v}}_{T}),\nabla\boldsymbol{w})_{T}$ $=(\nabla\boldsymbol{v}_{T},\nabla\boldsymbol{w})_{T}+\sum_{F\in\mathcal{F}_{T}}(\boldsymbol{v}_{F}-\boldsymbol{v}_{T},\nabla\boldsymbol{w}\cdot\boldsymbol{n}_{TF})_{F}\quad\forall\boldsymbol{w}\in[P^{k+1}(T)]^{2},$ (2.1)
@@
Approximation property of $\boldsymbol{r}_{T}^{k+1}$: There exists a real number $C>0$, depending on $\varrho$ but independent of $h_{T}$ such that, for all $\boldsymbol{v}\in[H^{s+1}(T)]^{2}$ for some $s\in\{0,1,\ldots,k+1\}$,
@@
For $s\in\{1,2,...,k+1\}$ and $\boldsymbol{v}\in[H^{s+1}(T)]^{2}$ we also have the approximation property
@@
Local advection reconstruction operator: $\boldsymbol{G}_{\boldsymbol{b},T}^{k}:\underline{\boldsymbol{U}}_{T}^{k}\rightarrow[P^{k}(T)]^{2}$ is defined as follows: For any $\underline{\boldsymbol{v}}_{T}\in\underline{\boldsymbol{U}}_{T}^{k}$
$(\boldsymbol{G}_{\boldsymbol{b},T}^{k}(\underline{\boldsymbol{v}}_{T}),\boldsymbol{w}_{T})_{T}=(\boldsymbol{b}\cdot\nabla\boldsymbol{v}_{T},\boldsymbol{w}_{T})_{T}+\sum_{F\in\mathcal{F}_{T}}(\boldsymbol{b}_{TF}(\boldsymbol{v}_{F}-\boldsymbol{v}_{T}),\boldsymbol{w}_{T})\quad\forall\boldsymbol{w}_{T}\in[P^{k}(T)]^{2}.$ (2.3)
@@ 2.2 A local Projection Setting
Let $\mathcal{M}_{h}$ be a finite decomposition of the domain $\Omega$ into open polygons, possibly overlapping so that $\cup_{M\in\mathcal{M}_{h}}\bar{M}=\bar{\Omega}$ and each $M\in\mathcal{M}_{h}$ is a collection of $T\in\mathcal{T}_{h}$. We assume that for all $M\in\mathcal{M}_{h}$ there exists a constant $C$ such that for any $M\in\mathcal{M}_{h}$ the cardinality of the set $\{N\in\mathcal{M}_{h}:N\cap M\neq\phi\}\leq C$. Let $h_{M}$ denote the diameter of the cell $M$. We also assume that for any polygon $T\in\mathcal{T}_{h}$ inside $M\in\mathcal{M}_{h}$, $h_{M}\lesssim h_{T}$. Let $K_{M}:[L^{2}(M)]^{2}\rightarrow[P^{k-1}(M)]^{2}$ be a bounded linear operator defined by $K_{M}:=Id_{M}-\pi^{k-1}_{M}$, $Id_{M}$ being the identity map.
Let $\boldsymbol{b}_{M}$ to be a piecewise constant approximation of $\boldsymbol{b}$ on $M$ such that $\left\|\boldsymbol{b}_{M}\right\|_{0,\infty,M}\leq C\left\|\boldsymbol{b}\right\|_{0,\infty,M}$ and $\left\|\boldsymbol{b}-\boldsymbol{b}_{M}\right\|_{0,\infty,M}\leq Ch_{M}\lvert\boldsymbol{b}\rvert_{1,\infty,M}$. For each cell $T$ contained in $M$, we define a local reconstruction $\boldsymbol{G}_{\boldsymbol{b}_{M},T}:\underline{\boldsymbol{U}}_{T}^{k}\rightarrow[P^{k}(T)]^{2}$ as follows: For $\underline{\boldsymbol{v}}_{T}\in\underline{\boldsymbol{U}}_{T}^{k}$
$(\boldsymbol{G}_{\boldsymbol{b}_{M},T}^{k}(\underline{\boldsymbol{v}}_{T}),\boldsymbol{w}_{T})_{T}=(\boldsymbol{b}_{M}\cdot\nabla\boldsymbol{v}_{T},\boldsymbol{w}_{T})_{T}+\sum_{F\in\mathcal{F}_{T}}(\boldsymbol{b}_{TF}(\boldsymbol{v}_{F}-\boldsymbol{v}_{T}),\boldsymbol{w}_{T})_{F}\quad\forall\boldsymbol{w}_{T}\in[P^{k}(T)]^{2}.$
Define $\boldsymbol{G}_{\boldsymbol{b}_{M},M}(\underline{\boldsymbol{v}}_{h})$ such that $\boldsymbol{G}_{\boldsymbol{b}_{M},M}(\underline{\boldsymbol{v}}_{h})\rvert_{T}=\boldsymbol{G}_{\boldsymbol{b}_{M},T}(\underline{\boldsymbol{v}}_{T})$ for each $T\subset M$. In this article, we propose the following local projection stabilisation $A_{S,h}:\underline{\boldsymbol{U}}_{h,0}^{k}\times\underline{\boldsymbol{U}}_{h,0}^{k}\to\mathbb{R}$ defined by
$A_{S,h}(\underline{\boldsymbol{v}}_{h},\underline{\boldsymbol{w}}_{h})=\sum_{M\in\mathcal{M}_{h}}\tau_{M}(K_{M}(\boldsymbol{G}_{\boldsymbol{b}_{M},M}(\underline{\boldsymbol{v}}_{h})),K_{M}(\boldsymbol{G}_{\boldsymbol{b}_{M},M}(\underline{\boldsymbol{w}}_{h})))_{M},$ (2.5)
where $\left\|\boldsymbol{b}\right\|_{0,\infty,M}\tau_{M}\sim h_{M}$. We obtain an estimate for $\left\|\boldsymbol{G}_{\boldsymbol{b},T}^{k}(\underline{\boldsymbol{u}}_{h})-\boldsymbol{G}_{\boldsymbol{b}_{M},T}^{k}(\underline{\boldsymbol{u}}_{h})\right\|$ as follows. Using the definition of the reconstructions $\boldsymbol{G}_{\boldsymbol{b},T}$ and $\boldsymbol{G}_{\boldsymbol{b}_{M},T}$ along with the approximation property of $\boldsymbol{b}_{M}$ we get
@@
$\leq Ch_{M}^{1/2}\left\|\boldsymbol{b}\right\|_{0,\infty}^{1/2}\left\|\boldsymbol{b}\right\|_{1,\infty}^{1/2}\left\|\nabla\boldsymbol{u}_{T}\right\|_{T}\left\|\boldsymbol{G}_{\boldsymbol{b},T}^{k}(\underline{\boldsymbol{u}}_{h})-\boldsymbol{G}_{\boldsymbol{b}_{M},T}^{k}(\underline{\boldsymbol{u}}_{h})\right\|_{T}.$
@@
$\left\|\boldsymbol{G}_{\boldsymbol{b},T}^{k}(\underline{\boldsymbol{u}}_{h})-\boldsymbol{G}_{\boldsymbol{b}_{M},T}^{k}(\underline{\boldsymbol{u}}_{h})\right\|_{T}\leq Ch_{M}^{1/2}\left\|\boldsymbol{b}\right\|_{0,\infty}^{1/2}\left\|\boldsymbol{b}\right\|_{1,\infty}^{1/2}\left\|\nabla\boldsymbol{u}_{T}\right\|_{T}.$ (2.6)
@@
Note that the decomposition $\mathcal{M}_{h}$ can be taken to be the original decomposition $\mathcal{T}_{h}$. The results proven in Sections 3 and 4 still hold with $\mathcal{M}_{h}=\mathcal{T}_{h}$. However, considering an overlapping decomposition $\mathcal{M}_{h}$ can significantly decrease the number of degrees of freedom required for the local projection and makes the local projection stabilisation more robust with respect to the choice of stabilisation parameter $\tau_{M}$, see MR2670000 .
@@ 2.3 Discrete Formulation
In this section, we introduce the discrete HHO-LPS scheme for the Oseen problem (0.1). The discrete problem is defined as follows: Find $(\underline{\boldsymbol{u}}_{h},p_{h})\in\underline{\boldsymbol{U}}_{h,0}^{k}\times P_{h}^{k}$ such that
@@
where, the combined bilinear form $A_{h}^{LP}(\cdot,\cdot)$ consists of the following parts:
@@
The viscosity bilinear form $A_{\epsilon,h}$: We use the local velocity reconstruction operator defined in (2.1) to define the viscosity term $A_{\epsilon,T}$. The local viscous bilinear form $A_{\epsilon,T}:\underline{\boldsymbol{U}}_{T}^{k}\times\underline{\boldsymbol{U}}_{T}^{k}\rightarrow\mathbb{R}$ is defined as
@@
$A_{N,h}(\underline{\boldsymbol{v}}_{h},\underline{\boldsymbol{w}}_{h}):=\sum_{T\in\mathcal{T}_{h}}\sum_{F\in\mathcal{F}_{T}}((\boldsymbol{v}_{T}-\boldsymbol{v}_{F})\cdot\boldsymbol{n},(\boldsymbol{w}_{T}-\boldsymbol{w}_{F})\cdot\boldsymbol{n})_{F}.$
@@ 3 Wellposedness of Discrete Formulation
$\left\|\underline{\boldsymbol{v}}_{h}\right\|^{2}_{\rm{dG}}:=\sum_{T\in\mathcal{T}_{h}}\Big{(}\left\|\nabla\boldsymbol{v}_{T}\right\|^{2}_{T}+\sum_{F\in\mathcal{F}_{T}}\frac{1}{h_{F}}\left\|\boldsymbol{v}_{F}-\boldsymbol{v}_{T}\right\|_{F}^{2}\Big{)},$
$\left\|\underline{\boldsymbol{v}}_{h}\right\|^{2}_{\epsilon,h}:=A_{\epsilon,h}(\underline{\boldsymbol{v}}_{h},\underline{\boldsymbol{v}}_{h}),\quad\left\|\underline{\boldsymbol{v}}_{h}\right\|^{2}_{\boldsymbol{b}}:=\sum_{T\in\mathcal{T}_{h}}\Big{(}\sigma\left\|\boldsymbol{v}_{T}\right\|^{2}_{T}+\sum_{F\in\mathcal{F}_{T}}\int_{F}\frac{\lvert\boldsymbol{b}_{TF}\rvert}{2}(\boldsymbol{v}_{F}-\boldsymbol{v}_{T})^{2}\Big{)}.$
In the proof of the stability of our discrete scheme, we will use the fact that the discontinuous Galerkin norm $\epsilon^{1/2}\left\|\cdot\right\|_{{\rm dG}}$ and the norm $\left\|\cdot\right\|_{\epsilon,h}$ are equivalent; see Piet_Jero_HHO_Book_20 .
@@
where $\omega=\max_{M\in\mathcal{M}_{h}}\frac{h_{M}^{2}}{\epsilon+\sigma h_{M}^{2}}$. We also assume that there exists a constant $\sigma_{0}$ such that $\sigma\leq\sigma_{0}$.
@@
The first and third terms of the above equation give
@@
For the second term of (3.2), we use the definition of $A_{\boldsymbol{b},T}$ in (2.3) and apply the integration by parts on $(\boldsymbol{b}\cdot\nabla\underline{\boldsymbol{u}}_{T},\underline{\boldsymbol{u}}_{T})_{T}$ along with the assumption ${\rm div~{}}\boldsymbol{b}=0$ to obtain
@@
$=\sum_{T\in\mathcal{T}_{h}}\Big{(}\sigma\left\|\boldsymbol{u}_{T}\right\|^{2}_{T}+\sum_{F\in\mathcal{F}_{T}}\frac{\lvert\boldsymbol{b}_{TF}\rvert}{2}\left\|\boldsymbol{u}_{F}-\boldsymbol{u}_{T}\right\|^{2}_{F}\Big{)}=\left\|\underline{\boldsymbol{u}}_{h}\right\|_{\boldsymbol{b}}^{2}.$ (3.4)
@@
for some positive constant $C$ independent of $h$ and $\epsilon$.
@@
For given $(\underline{\boldsymbol{u}}_{h},p_{h})\in\underline{\boldsymbol{U}}_{h,0}^{k}\times P_{h}^{k}$, let $\boldsymbol{x}_{M}=\boldsymbol{G}_{\boldsymbol{b},M}^{k}(\underline{\boldsymbol{u}}_{T})+\nabla_{h}p_{h}\rvert_{M}$ for $M\in\mathcal{M}_{h}$. We define $\boldsymbol{z}_{M}=\gamma_{M}\pi_{M}^{k-1}\boldsymbol{x}_{M}$ on $M$ and extend to $\Omega$ by zero. Using the boundedness of $\pi_{M}^{k-1}$ we have
$\left\|\boldsymbol{z}_{M}\right\|_{M}$ $\leq C\gamma_{M}\left\|\boldsymbol{x}_{M}\right\|_{M}.$ (3.6)
Now define the global function $\underline{\boldsymbol{z}}_{h}=((\boldsymbol{z}_{T})_{T\in\mathcal{T}_{h}},(\boldsymbol{z}_{F})_{F\in\mathcal{F}_{h}})$, where $\boldsymbol{z}_{T}=\sum_{M}\boldsymbol{z}_{M}|_{T}$ and $z_{F}=\boldsymbol{0}$ for all $F\in\mathcal{F}_{h}$. From the properties of the decomposition $\mathcal{M}_{h}$ and using the inverse inequality, the trace inequality and the equivalence $h_{M}\lesssim h_{F}$ we have the following bounds
@@
$\sum_{T\in\mathcal{T}_{h}}\left\|\nabla\boldsymbol{z}_{T}\right\|^{2}_{T}\leq C\sum_{M\in\mathcal{M}_{h}}\frac{1}{h_{M}^{2}}\left\|\boldsymbol{z}_{M}\right\|^{2}_{M},\quad\sum_{T\in\mathcal{T}}\sum_{F\in\mathcal{F}_{T}}\frac{1}{h_{F}}\left\|\boldsymbol{z}_{T}\right\|^{2}_{F}\leq C\sum_{M\in\mathcal{M}_{h}}\frac{1}{h_{M}^{2}}\left\|\boldsymbol{z}_{M}\right\|^{2}_{M}.$ (3.7)

### v2

@@ 2 Discrete Oseen Problem with LPS Stabilisation
In this section, we introduce the LPS stabilised discrete formulation for the Oseen problem (1.2) on the hybrid space $\underline{\boldsymbol{U}}_{h,0}^{k}\times P_{h}^{k}$. This section is divided into three sub-sections. The first part defines some reconstruction operators that are essential to define the HHO method. In the second part, we discuss a generalised local projection stabilisation setup. The discrete LPS-HHO method is defined in the third subsection.
@@ 2.1 Local Reconstructions
We define three reconstruction operators on the local spaces $\underline{\boldsymbol{U}}_{T}^{k}$, see [1]. These are used to define the discrete HHO bilinear form in (2.7).
Local velocity reconstruction: The velocity reconstruction operator $\boldsymbol{r}_{T}^{k+1}:\underline{\boldsymbol{U}}_{T}^{k}\rightarrow[P^{k+1}(T)]^{d}$ is defined as follows: For any $\underline{\boldsymbol{v}}_{T}\in\underline{\boldsymbol{U}}_{T}^{k}$, $\boldsymbol{r}_{T}^{k+1}\underline{\boldsymbol{v}}_{T}$ must satisfy
$(\nabla(\boldsymbol{r}_{T}^{k+1}\underline{\boldsymbol{v}}_{T}),\nabla\boldsymbol{w})_{T}$ $=(\nabla\boldsymbol{v}_{T},\nabla\boldsymbol{w})_{T}+\sum_{F\in\mathcal{F}_{T}}(\boldsymbol{v}_{F}-\boldsymbol{v}_{T},\nabla\boldsymbol{w}\boldsymbol{n}_{TF})_{F}\quad\forall\boldsymbol{w}\in[P^{k+1}(T)]^{d},$ (2.1)
@@
Approximation property of $\boldsymbol{r}_{T}^{k+1}$: There exists a real number $C>0$, depending on $\varrho$ but independent of $h_{T}$ such that, for all $\boldsymbol{v}\in[H^{s+1}(T)]^{d}$ for some $s\in\{0,1,\ldots,k+1\}$,
@@
For $s\in\{1,2,...,k+1\}$ and $\boldsymbol{v}\in[H^{s+1}(T)]^{d}$ we also have the approximation property
@@
Local advection reconstruction operator: $\boldsymbol{G}_{\boldsymbol{b},T}^{k}:\underline{\boldsymbol{U}}_{T}^{k}\rightarrow[P^{k}(T)]^{d}$ is defined as follows: For any $\underline{\boldsymbol{v}}_{T}\in\underline{\boldsymbol{U}}_{T}^{k}$
$(\boldsymbol{G}_{\boldsymbol{b},T}^{k}(\underline{\boldsymbol{v}}_{T}),\boldsymbol{w}_{T})_{T}=(\boldsymbol{b}\cdot\nabla\boldsymbol{v}_{T},\boldsymbol{w}_{T})_{T}+\sum_{F\in\mathcal{F}_{T}}(\boldsymbol{b}_{TF}(\boldsymbol{v}_{F}-\boldsymbol{v}_{T}),\boldsymbol{w}_{T})_{F}\quad\forall\boldsymbol{w}_{T}\in[P^{k}(T)]^{d}.$ (2.3)
@@ 2.2 A Local Projection Setting
Let $\mathcal{M}_{h}$ be a finite decomposition of the domain $\Omega$ into open subsets, possibly overlapping so that $\cup_{M\in\mathcal{M}_{h}}\bar{M}=\bar{\Omega}$ and each $M\in\mathcal{M}_{h}$ is a collection of $T\in\mathcal{T}_{h}$. We assume that there exists a constant $C$ such that for any $M\in\mathcal{M}_{h}$ the cardinality of the set $\{N\in\mathcal{M}_{h}:N\cap M\neq\phi\}\leq C$. Let $h_{M}$ denote the diameter of the cell $M$. We also assume that for any cell $T\in\mathcal{T}_{h}$ inside $M\in\mathcal{M}_{h}$, $h_{M}\lesssim h_{T}$. Let $K_{M}:[L^{2}(M)]^{d}\rightarrow[P^{k-1}(M)]^{d}$ be a bounded linear operator defined by $K_{M}:=Id_{M}-\pi^{k-1}_{M}$, $Id_{M}$ being the identity map.
Let $\boldsymbol{b}_{M}$ to be a piecewise constant approximation of $\boldsymbol{b}$ on $M$ such that $\left\|\boldsymbol{b}_{M}\right\|_{0,\infty,M}\leq C\left\|\boldsymbol{b}\right\|_{0,\infty,M}$ and $\left\|\boldsymbol{b}-\boldsymbol{b}_{M}\right\|_{0,\infty,M}\leq Ch_{M}\lvert\boldsymbol{b}\rvert_{1,\infty,M}$. For each cell $T$ contained in $M$, we define a local reconstruction $\boldsymbol{G}_{\boldsymbol{b}_{M},T}^{k}:\underline{\boldsymbol{U}}_{T}^{k}\rightarrow[P^{k}(T)]^{d}$ as follows: For $\underline{\boldsymbol{v}}_{T}\in\underline{\boldsymbol{U}}_{T}^{k}$
$(\boldsymbol{G}_{\boldsymbol{b}_{M},T}^{k}(\underline{\boldsymbol{v}}_{T}),\boldsymbol{w}_{T})_{T}=(\boldsymbol{b}_{M}\cdot\nabla\boldsymbol{v}_{T},\boldsymbol{w}_{T})_{T}+\sum_{F\in\mathcal{F}_{T}}(\boldsymbol{b}_{TF}(\boldsymbol{v}_{F}-\boldsymbol{v}_{T}),\boldsymbol{w}_{T})_{F}\quad\forall\boldsymbol{w}_{T}\in[P^{k}(T)]^{d}.$
Define $\boldsymbol{G}_{\boldsymbol{b}_{M},M}^{k}(\underline{\boldsymbol{v}}_{h})$ such that $\boldsymbol{G}_{\boldsymbol{b}_{M},M}^{k}(\underline{\boldsymbol{v}}_{h})\rvert_{T}=\boldsymbol{G}_{\boldsymbol{b}_{M},T}^{k}(\underline{\boldsymbol{v}}_{T})$ for each $T\subset M$ and $\boldsymbol{G}_{\boldsymbol{b},M}^{k}(\underline{\boldsymbol{v}}_{h})$ as $\boldsymbol{G}_{\boldsymbol{b},M}^{k}(\underline{\boldsymbol{v}}_{h})|_{T}=\boldsymbol{G}_{\boldsymbol{b},T}^{k}(\underline{\boldsymbol{v}}_{T})$. In this article, we propose the following local projection stabilisation $A_{S,h}:\underline{\boldsymbol{U}}_{h,0}^{k}\times\underline{\boldsymbol{U}}_{h,0}^{k}\to\mathbb{R}$ defined by
$A_{S,h}(\underline{\boldsymbol{v}}_{h},\underline{\boldsymbol{w}}_{h})=\sum_{M\in\mathcal{M}_{h}}\tau_{M}(K_{M}(\boldsymbol{G}_{\boldsymbol{b}_{M},M}^{k}(\underline{\boldsymbol{v}}_{h})),K_{M}(\boldsymbol{G}_{\boldsymbol{b}_{M},M}^{k}(\underline{\boldsymbol{w}}_{h})))_{M},$ (2.5)
where $\left\|\boldsymbol{b}\right\|_{0,\infty,M}\tau_{M}\leq Ch_{M}$ with the constant $C$ independent of $M,F,h$ and the data of (0.1), see [46, eqn. (2.5)]. We obtain an estimate for $\left\|\boldsymbol{G}_{\boldsymbol{b},T}^{k}(\underline{\boldsymbol{u}}_{h})-\boldsymbol{G}_{\boldsymbol{b}_{M},T}^{k}(\underline{\boldsymbol{u}}_{h})\right\|$ as follows. Using the definition of the reconstructions $\boldsymbol{G}_{\boldsymbol{b},T}^{k}$ and $\boldsymbol{G}_{\boldsymbol{b}_{M},T}^{k}$ along with the approximation property of $\boldsymbol{b}_{M}$ we get
@@
$\leq Ch_{M}^{1/2}\left\|\boldsymbol{b}\right\|_{0,\infty,M}^{1/2}\left\|\boldsymbol{b}\right\|_{1,\infty,M}^{1/2}\left\|\nabla\boldsymbol{u}_{T}\right\|_{T}\left\|\boldsymbol{G}_{\boldsymbol{b},T}^{k}(\underline{\boldsymbol{u}}_{h})-\boldsymbol{G}_{\boldsymbol{b}_{M},T}^{k}(\underline{\boldsymbol{u}}_{h})\right\|_{T}.$
@@
$\left\|\boldsymbol{G}_{\boldsymbol{b},T}^{k}(\underline{\boldsymbol{u}}_{h})-\boldsymbol{G}_{\boldsymbol{b}_{M},T}^{k}(\underline{\boldsymbol{u}}_{h})\right\|_{T}\leq Ch_{M}^{1/2}\left\|\boldsymbol{b}\right\|_{0,\infty,M}^{1/2}\left\|\boldsymbol{b}\right\|_{1,\infty,M}^{1/2}\left\|\nabla\boldsymbol{u}_{T}\right\|_{T}.$ (2.6)
@@
Note that the decomposition $\mathcal{M}_{h}$ can be taken to be the original decomposition $\mathcal{T}_{h}$. The results proven in Sections 3 and 4 still hold with $\mathcal{M}_{h}=\mathcal{T}_{h}$. However, considering an overlapping decomposition $\mathcal{M}_{h}$ can significantly decrease the number of degrees of freedom required for the local projection and makes the local projection stabilisation more robust with respect to the choice of stabilisation parameter $\tau_{M}$, see [45].
@@ 2.3 Discrete Formulation
In this section, we introduce the discrete HHO-LPS method for the Oseen problem (0.1). The discrete problem is defined as follows: Find $(\underline{\boldsymbol{u}}_{h},p_{h})\in\underline{\boldsymbol{U}}_{h,0}^{k}\times P_{h}^{k}$ such that
@@
where the combined bilinear form $A_{h}^{LP}(\cdot,\cdot)$ consists of the following parts:
@@
The viscosity bilinear form $A_{\epsilon,h}$: We use the local velocity reconstruction operator defined in (2.1) to define the viscosity term $A_{\epsilon,h}$. The local viscous bilinear form $A_{\epsilon,T}:\underline{\boldsymbol{U}}_{T}^{k}\times\underline{\boldsymbol{U}}_{T}^{k}\rightarrow\mathbb{R}$ is defined as
@@
$A_{N,h}(\underline{\boldsymbol{v}}_{h},\underline{\boldsymbol{w}}_{h}):=\sum_{T\in\mathcal{T}_{h}}\sum_{F\in\mathcal{F}_{T}}((\boldsymbol{v}_{T}-\boldsymbol{v}_{F})\cdot\boldsymbol{n}_{TF},(\boldsymbol{w}_{T}-\boldsymbol{w}_{F})\cdot\boldsymbol{n}_{TF})_{F}.$
@@ 3 Wellposedness of Discrete Formulation
$\left\|\underline{\boldsymbol{v}}_{h}\right\|^{2}_{{1,h}}:=\sum_{T\in\mathcal{T}_{h}}\Big{(}\left\|\nabla\boldsymbol{v}_{T}\right\|^{2}_{T}+\sum_{F\in\mathcal{F}_{T}}\frac{1}{h_{F}}\left\|\boldsymbol{v}_{F}-\boldsymbol{v}_{T}\right\|_{F}^{2}\Big{)},$
$\left\|\underline{\boldsymbol{v}}_{h}\right\|^{2}_{\epsilon,h}:=A_{\epsilon,h}(\underline{\boldsymbol{v}}_{h},\underline{\boldsymbol{v}}_{h}),\quad\left\|\underline{\boldsymbol{v}}_{h}\right\|^{2}_{\boldsymbol{b}}:=\sum_{T\in\mathcal{T}_{h}}\Big{(}\sigma\left\|\boldsymbol{v}_{T}\right\|^{2}_{T}+\sum_{F\in\mathcal{F}_{T}}\int_{F}\frac{\lvert\boldsymbol{b}_{TF}\rvert}{2}(\boldsymbol{v}_{F}-\boldsymbol{v}_{T})\cdot(\boldsymbol{v}_{F}-\boldsymbol{v}_{T})\Big{)}.$
In the proof of the stability of our discrete scheme, we will use the fact that the norms $\epsilon^{1/2}\left\|\cdot\right\|_{1,h}$ and $\left\|\cdot\right\|_{\epsilon,h}$ are equivalent; see [25].
@@
where $\omega=\max_{M\in\mathcal{M}_{h}}\frac{h_{M}^{2}\left\|\boldsymbol{b}\right\|_{1,\infty,M}}{\epsilon+\sigma h_{M}^{2}}$.
@@
The first and third terms of the above equation read as
@@
For the second term of (3.2), we use the definition of $A_{\boldsymbol{b},T}$ in (2.3) and apply the integration by parts along with the assumption ${\rm div~{}}\boldsymbol{b}=0$ to obtain (see [1, eqn. (19)] for more details)
@@
$=\sum_{T\in\mathcal{T}_{h}}\Big{(}\sigma\left\|\boldsymbol{u}_{T}\right\|^{2}_{T}+\sum_{F\in\mathcal{F}_{T}}\frac{1}{2}\left\|\lvert\boldsymbol{b}_{TF}\rvert^{1/2}(\boldsymbol{u}_{F}-\boldsymbol{u}_{T})\right\|^{2}_{F}\Big{)}=\left\|\underline{\boldsymbol{u}}_{h}\right\|_{\boldsymbol{b}}^{2}.$ (3.4)
@@
for some positive constant $C_{1}$ independent of $h$ and $\epsilon$.
@@
We follow some steps of [46, Theorem 4.1] for the proof of this lemma. For given $(\underline{\boldsymbol{u}}_{h},p_{h})\in\underline{\boldsymbol{U}}_{h,0}^{k}\times P_{h}^{k}$, let $\boldsymbol{x}_{M}=\boldsymbol{G}_{\boldsymbol{b},M}^{k}(\underline{\boldsymbol{u}}_{h})+\nabla_{h}p_{h}$ for $M\in\mathcal{M}_{h}$. We define $\boldsymbol{z}_{M}=\gamma_{M}\pi_{M}^{k-1}\boldsymbol{x}_{M}$ on $M$ and extend to $\Omega$ by zero. Using the boundedness of $\pi_{M}^{k-1}$ we have
$\left\|\boldsymbol{z}_{M}\right\|_{M}$ $\leq\gamma_{M}\left\|\boldsymbol{x}_{M}\right\|_{M}.$ (3.6)
Now define the global function $\underline{\boldsymbol{z}}_{h}=((\boldsymbol{z}_{T})_{T\in\mathcal{T}_{h}},(\boldsymbol{z}_{F})_{F\in\mathcal{F}_{h}})$, where $\boldsymbol{z}_{T}=\sum_{M\supset T}\boldsymbol{z}_{M}|_{T}$ and $z_{F}=\boldsymbol{0}$ for all $F\in\mathcal{F}_{h}$. From the properties of the decomposition $\mathcal{M}_{h}$ and using the inverse inequality, the trace inequality, and the equivalence $h_{M}\lesssim h_{F}$, we have the following bounds:
@@
$\sum_{T\in\mathcal{T}_{h}}\left\|\nabla\boldsymbol{z}_{T}\right\|^{2}_{T}\leq C\sum_{M\in\mathcal{M}_{h}}\frac{1}{h_{M}^{2}}\left\|\boldsymbol{z}_{M}\right\|^{2}_{M},\quad\sum_{T\in\mathcal{T}_{h}}\sum_{F\in\mathcal{F}_{T}}\frac{1}{h_{F}}\left\|\boldsymbol{z}_{T}\right\|^{2}_{F}\leq C\sum_{M\in\mathcal{M}_{h}}\frac{1}{h_{M}^{2}}\left\|\boldsymbol{z}_{M}\right\|^{2}_{M}.$ (3.7)